If a son is 1/4 his mother's age, and she is 21 years older than her son, how old is the son?
Answer by blueplanet(7)

Let the mother's age be X.
Her son age is X-21
According to the equation, the son is 1/4 his mother's age. This means that:
X-21 = X/4 ---- (1)
Solving this equation, we get X=28, which is the mother's age. The son's age is 28-21= 7 years old.